# SAM 2: Segment Anything in Images and Videos

### v1

@@ G.2 Dataset card for SA-V dataset
    Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation. See detail in 6.1.3.
@@
G.3 Data annotation card
@@
    Consider the intended context of use of the dataset and the individuals and communities that may be impacted by a model trained on this dataset. Are these communities represented in your annotator pool? The SA-V dataset is a geographically diverse, publicly available, video segmentation dataset, as discussed in 5.2. In addition, we analyze the responsible AI axes of a model trained on the dataset, as discussed in 6.1.3.

### v2

@@ H.2 Dataset card for SA-V dataset
    Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation. See detail in E.1.1.
@@
H.3 Data annotation card
@@
    Consider the intended context of use of the dataset and the individuals and communities that may be impacted by a model trained on this dataset. Are these communities represented in your annotator pool? The SA-V dataset is a geographically diverse, publicly available, video segmentation dataset, as discussed in 5.2. In addition, we analyze the responsible AI axes of a model trained on the dataset, as discussed in E.1.1.